# Bias-Reduced Estimation of Structural Equation Models

### v1

@@ Abstract
Finite-sample bias is a pervasive challenge in the estimation of structural equation models (SEMs), especially when sample sizes are small or measurement reliability is low. A range of methods have been proposed to improve finite-sample bias in the SEM literature, ranging from analytic bias corrections to resampling-based techniques, with each carrying trade-offs in scope, computational burden, and statistical performance. We apply the reduced-bias M-estimation framework [[, RBM,]J. R. Stat. Soc. Series B Stat. Methodol.]kosmidis2024empirical to SEMs. The RBM framework is attractive as it requires only first- and second-order derivatives of the log-likelihood, which renders it both straightforward to implement, and computationally more efficient compared to resampling-based alternatives such as bootstrap and jackknife. It is also robust to departures from modelling assumptions. Through extensive simulations studies under a range of experimental conditions, we illustrate that RBM estimators consistently reduce mean bias in the estimation of SEMs without inflating mean squared error. They also deliver improvements in both median bias and inference relative to maximum likelihood estimators, while maintaining robustness under non-normality. Our findings suggest that RBM offers a promising, practical, and broadly applicable tool for mitigating bias in the estimation of SEMs, particularly in small-sample research contexts.
@@ 1 Introduction
Structural Equation Models (SEMs) are widely used statistical models that allow for the specification, estimation, and testing of complex relationships among observed and latent variables. Despite their versatility and theoretical appeal, the practical use of SEMs in empirical research is often challenged by small sample sizes. Small samples are particularly common in fields where data collection is expensive, time-consuming, or logistically constrained, such as clinical psychology, developmental research, and neuroscience [(49)]. Small samples are also often an inevitable consequence of studying rare populations. For example, research involving children with burned facial injuries is inherently limited by the small number of individuals who meet these specific criteria [(41)]. Similarly, studies focusing on rare genetic disorders [(17)], elite athletes [(23), (16)], or unique cultural groups [(32)] often face strict limitations in sample size due to the rarity of the target population. In such contexts, the use of SEMs remains desirable for testing complex theoretical models, but the challenges posed by small samples become unavoidable and cannot be ignored.
@@
Motivated by these appealing properties, the goal of this paper is to investigate the performance of the RBM estimation framework for SEM. The RBM method is straightforward to implement for any SEM, and we provide accompanying R code in the supplementary materials to facilitate its application. This work builds on [(12)], who investigated resampling-based approaches to bias reduction for two types of SEMs: a two-factor SEM, and a latent growth curve model.
The remainder of the paper is structured as follows. We begin by briefly outlining the SEM framework and introducing the notation used throughout the paper. We then present a concise overview of several bias-reduction techniques, including the method proposed by [(37)], the jackknife and bootstrap approaches, and the recently introduced RBM method. Following this, we report the results of a comprehensive simulation study designed to evaluate the performance of the RBM approach in comparison to existing alternatives. Finally, we conclude with a discussion of the findings and offer practical recommendations for applied researchers.
@@ 4.1 Models
$\Lambda=\begin{bmatrix}1&0\\ 1&1\\ \vdots&\vdots\\ 1&9\\ \end{bmatrix}\quad\text{and}\quad B=\begin{bmatrix}0&0\\ 0&0\\ \end{bmatrix}$.
@@ 4.4 Computational Note
All analyses were conducted in R [(42)]. ML estimation, including robust standard errors, was performed using the {lavaan} package [(43), with se = "robust.huber.white",]. The eRBM and iRBM estimators and robust standard errors were implemented using our freely available {brlavaan} package, which also provides R functions for the [(37)] correction, bootstrap, and jackknife procedures tailored to the two-factor SEM and latent GCM considered here. The R package {lme4} [(1)] was used for fitting the latent GCM as a LMM. All code used to perform the simulations, along with the computed simulation results, is openly available on the Open Science Framework repository via the URL provided in the supplementary materials.
@@ 6 Discussion
Compared to traditional resampling-based approaches such as jackknife and bootstrap, iRBM and eRBM achieved a more favourable bias–variance trade-off. While resampling sometimes produced slightly lower bias, this was often accompanied by greater variability in the estimates, and in the case of bootstrap, a much higher computational burden. In contrast, iRBM and eRBM deliver bias reduction at a fraction of the computational cost, requiring only a single optimisation rather than repeated refitting across hundreds of resamples. For a modest sample size of $n=100$, eRBM required only $0.605$ seconds per replication on average for the two-factor SEM (c.f. bootstrap $53.2$ seconds; jackknife $9.77$ seconds), and $0.317$ seconds for the latent GCM (c.f. bootstrap $42.8$ seconds; jackknife $8.56$ seconds), representing significant reductions in runtime compared to resampling methods. Although iRBM was slower than eRBM due to the penalized likelihood optimisation (averaging $13.0$ seconds per replication for the two-factor SEM and $1.20$ seconds for the latent GCM), it was still substantially faster than bootstrap while yielding comparable or superior statistical performance. REML, while theoretically optimal under normality, performed poorly under non-normal data, underscoring the importance of robust approaches such as RBM that do not rely heavily on distributional assumptions.
@@
From a computational perspective, our current implementation is intentionally simple, relying on numerical derivatives and basic optimisation routines. For iRBM, each evaluation of the penalized log-likelihood requires computing the inverse Hessian and the cross-product of the score contributions at the current $\vartheta$, both of which are obtained numerically. This repeated numerical differentiation at every optimisation step is the primary contributor to the higher runtime of iRBM. In contrast, eRBM performs the bias correction post hoc, which avoids repeated evaluations inside the optimiser and is therefore somewhat faster for SEMs. eRBM would be useful in settings where a model has already been fitted and a quick, post-hoc adjustment is desired. Nevertheless, our recommendation—reflected in the default of the {brlavaan} package—is to use iRBM, as it performs the correction implicitly and is less prone to producing infeasible estimates. Future computational improvements could include analytic derivations of the score contributions and the $A(\vartheta)$ matrix for common SEM specifications, as well as more efficient matrix factorization and quasi-Newton updates to reduce the cost of recomputing the inverse Hessian. Such improvements could dramatically reduce runtimes, making iRBM competitive with eRBM while preserving its robustness advantages.
@@
Another direction for future work is to explore plugin penalties (e.g. sterzinger2023maximum, 98) to keep iRBM estimates within the interior of the parameter space. Such penalties could improve numerical stability, limit exploration of ill-conditioned regions, and potentially shorten runtimes. Although the proportion of unacceptable estimates was small, we did observe occasional iRBM estimation failures for $n\leq 50$ in both the two-factor SEM and latent GCM (Table 2). Light regularization of this type could therefore further enhance the stability of iRBM in these challenging small-sample scenarios where the likelihood surface is nearly flat. Extensions to more complex SEMs—such as mediation models, models with latent interactions, and those estimated using alternatives to ML (e.g., WLS, DWLS)—are also a natural next step, given the generality of the M-estimation framework underlying RBM. Further investigation into non-normal likelihoods and robust estimation approaches would broaden the reach of RBM-based bias reduction to a wider class of psychometric and structural modelling problems.
@@ Appendix A Derivatives
is referred to as the structure matrix of $\mathbf{M}$ (petersen2012matrix, 89, Eq. 134,). Here, $\mathbf{J}^{ij}\in\mathbb{R}^{p\times q}$ is the single-entry matrix with a ‘$1$’ at the $(i,j)$ position and 0 elsewhere. Write
@@
using the fact that $d\tilde{\mathbf{B}}=\tilde{\mathbf{B}}(d\mathbf{B})\tilde{\mathbf{B}}$ (petersen2012matrix, 89, Eq. 59). Furthermore,

### v2

@@ Abstract
Finite-sample bias is a pervasive challenge in the estimation of structural equation models (SEMs), especially when sample sizes are small or measurement reliability is low. A range of methods have been proposed to improve finite-sample bias in the SEM literature, ranging from analytic bias corrections to resampling-based techniques, with each carrying trade-offs in scope, computational burden, and statistical performance. We introduce the reduced-bias M-estimation framework [[, RBM,]J. R. Stat. Soc. Series B Stat. Methodol.]kosmidis2024empirical to SEMs. The RBM framework is attractive as it requires only first- and second-order derivatives of the log-likelihood, which renders it both straightforward to implement, and computationally more efficient compared to resampling-based alternatives such as bootstrap and jackknife. It is also robust to departures from modelling assumptions. Using the same simulation setup as in [(12)], we illustrate that RBM estimators consistently reduce mean bias in the estimation of SEMs without inflating mean squared error. They also deliver improvements in both median bias and inference relative to maximum likelihood estimators, while maintaining robustness under non-normality. Our findings suggest that RBM offers a promising, practical, and broadly applicable tool for mitigating bias in the estimation of SEMs, particularly in small-sample research contexts.
@@ 1 Introduction
Structural Equation Models (SEMs) are widely used statistical models that allow for the specification, estimation, and testing of complex relationships among observed and latent variables [(25), (52)]. Despite their versatility and theoretical appeal, the practical use of SEMs in empirical research is often challenged by small sample sizes. Small samples are particularly common in fields where data collection is expensive, time-consuming, or logistically constrained, such as clinical psychology, developmental research, and neuroscience [(55)]. Small samples are also often an inevitable consequence of studying rare populations. For example, research involving children with burned facial injuries is inherently limited by the small number of individuals who meet these specific criteria [(45)]. Similarly, studies focusing on rare genetic disorders [(17)], elite athletes [(24), (16)], or unique cultural groups [(36)] often face strict limitations in sample size due to the rarity of the target population. In such contexts, the use of SEMs remains desirable for testing complex theoretical models, but the challenges posed by small samples become unavoidable and cannot be ignored.
@@
Motivated by these appealing properties, the goal of this paper is to translate the RBM technology to the SEM framework and to investigate its performance by extending the simulation setup of [(12)], who investigated resampling-based approaches to bias reduction for two types of SEMs: a two-factor SEM, and a latent growth curve model. In addition, we provide accompanying R code in the supplementary materials to facilitate its application.
The remainder of the paper is structured as follows. We begin by briefly outlining the SEM framework and introducing the notation used throughout the paper in section 2. In section 3, we then present a concise overview of several bias-reduction techniques, including the method proposed by [(41)], the jackknife and bootstrap approaches, and the recently introduced RBM method. Following this, in sections 4 and 5, we report the results of two simulation studies designed to evaluate the performance of the RBM approach in comparison to a selection of alternatives. Finally, in section 6, we conclude with a discussion of the findings and offer practical recommendations for applied researchers.
@@ 3.4 Reduced-Bias M-Estimation
The practical implementation of RBM within the normal SEM framework requires translating the general quantities in (9) into expressions under the multivariate normal likelihood with model-implied mean vector $\boldsymbol{\mu}(\vartheta)$ and covariance matrix $\boldsymbol{\Sigma}(\vartheta)$ given by (3) and (4). Thus, the penalty components depend intricately on the derivatives of $\boldsymbol{\mu}(\vartheta)$ and $\boldsymbol{\Sigma}(\vartheta)$ with respect to the free parameters $\vartheta$ of the SEM. We derive analytic closed-form expressions of these first-order derivatives via the chain rule for a general SEM in Appendix A. Higher-order derivatives required to evaluate $\nabla P(\vartheta)$ are obtained by numerical differentiation of these first-order quantities, as is common in SEM software such as {lavaan} [(47)]. This strategy facilitates a robust application to a wide array of SEM specifications.
@@ 4.1 Models
$\boldsymbol{\Lambda}=\begin{bmatrix}1&0\\ 1&1\\ \vdots&\vdots\\ 1&9\\ \end{bmatrix}\quad\text{and}\quad\mathbf{B}=\begin{bmatrix}0&0\\ 0&0\\ \end{bmatrix}$.
@@ 4.4 Computational Note
All analyses were conducted in R [(46)]. Computations were performed on a 15-inch MacBook Air (2023) equipped with an Apple M2 chip (8-core CPU, 4 performance cores up to 3.50 GHz) and 16 GB of unified memory.
ML estimation, including robust standard errors, was performed using the {lavaan} package [(47), with se = "robust.huber.white",]. The eRBM and iRBM estimators and robust standard errors were implemented using our freely available {brlavaan} package, which also provides R functions for the [(41)] correction, bootstrap, and jackknife procedures tailored to the two-factor SEM and latent GCM considered here. The R package {lme4} [(1)] was used for fitting the latent GCM as a LMM. All code used to perform the simulations, along with the computed simulation results, is openly available on the Open Science Framework repository via the URL provided in the supplementary materials.
@@ 6 Discussion
Compared to traditional resampling-based approaches such as jackknife and bootstrap, iRBM and eRBM achieved a more favourable bias–variance trade-off. While resampling sometimes produced slightly lower bias, this was often accompanied by greater variability in the estimates, and in the case of bootstrap, a much higher computational burden. In contrast, iRBM and eRBM deliver bias reduction at a fraction of the computational cost, requiring only a single optimisation rather than repeated refitting across hundreds of resamples. For a modest sample size of $n=100$, eRBM required only $0.605$ seconds per replication on average for the two-factor SEM (c.f. bootstrap $53.2$ seconds; jackknife $9.77$ seconds), and $0.317$ seconds for the latent GCM (c.f. bootstrap $42.8$ seconds; jackknife $8.56$ seconds), representing significant reductions in runtime [^1] Runtimes were measured on a MacBook Air M2 (8-core CPU @ 3.50 GHz, 16 GB RAM). compared to resampling methods. Although iRBM was slower than eRBM due to the penalized likelihood optimisation (averaging $13.0$ seconds per replication for the two-factor SEM and $1.20$ seconds for the latent GCM), it was still substantially faster than bootstrap while yielding comparable or superior statistical performance. REML, while theoretically optimal under normality, performed poorly under non-normal data, underscoring the importance of robust approaches such as RBM that do not rely heavily on distributional assumptions.
@@
From a computational perspective, our implementation uses analytic casewise score vectors (Appendix A), while the Hessian and its inverse are obtained numerically at the current $\vartheta$. For iRBM, each evaluation of the penalized log-likelihood requires computing the inverse Hessian and the cross-product of the score contributions at the current $\vartheta$. This repeated numerical differentiation at every optimisation step is the primary contributor to the higher runtime of iRBM. By comparison, eRBM performs the bias correction post hoc, which avoids repeated evaluations inside the optimiser and is therefore somewhat faster for SEMs. eRBM would be useful in settings where a model has already been fitted and a quick, post-hoc adjustment is desired. Nevertheless, our recommendation—reflected in the default of the {brlavaan} package—is to use iRBM, as it performs the correction implicitly and is less prone to producing infeasible estimates. Future computational improvements could include analytic derivations of the $A(\vartheta)$ matrix for common SEM specifications, as well as more efficient matrix factorization and quasi-Newton updates to reduce the cost of recomputing the inverse Hessian. Such improvements could dramatically reduce runtimes, making iRBM competitive with eRBM while preserving its robustness advantages.
Relatedly, a computational trade-off presents itself as models become more complex. For instance, when more than two factors are involved and the dimension of the parameter vector $\vartheta$ expands significantly. Introducing additional latent variables necessitates estimating more factor loadings and a larger latent covariance matrix, which increases the size of the Hessian and variability matrix. Because of the aforementioned numerical routine involved, the runtime for iRBM is expected to increase non-linearly with the number of parameters. Then, the statistical arguments for using iRBM become even stronger in these high-dimensional settings. It is well-documented that as model complexity grows relative to sample size, the finite-sample properties of ML estimators deteriorate jackson2003revisiting (84, 114). This ‘model size effect’ typically manifests as increased bias in variance components smid2020bayesian (110) and factor correlations gerbing1985effects (77), particularly when measurement quality is low or loadings are weakly identified maccallum1999sample (92). Furthermore, the standard ML estimator becomes increasingly prone to non-convergence or inadmissible solutions (e.g., Heywood cases), which can compromise the feasibility of the post-hoc eRBM correction in practice. On the other hand, iRBM operates on a penalized likelihood surface that often exhibits improved numerical behaviour and mitigates instability. Consequently, while iRBM may require more computational time per replication in complex models, it effectively serves as a self-regularizing procedure that is less dependent on the existence of a stable initial ML solution, supporting our recommendation of iRBM as the preferred option for small-sample SEM.
@@
Returning to potential future work, another direction is to explore plugin penalties (e.g. sterzinger2023maximum, 111) to keep iRBM estimates within the interior of the parameter space. Such penalties could improve numerical stability, limit exploration of ill-conditioned regions, and potentially shorten runtimes. Although the proportion of unacceptable estimates was small, we did observe occasional iRBM estimation failures for $n\leq 50$ in both the two-factor SEM and latent GCM (Table 2). Light regularization of this type could therefore further enhance the stability of iRBM in these challenging small-sample scenarios where the likelihood surface is nearly flat. Extensions to more complex SEMs—such as mediation models, models with latent interactions, and those estimated using alternatives to ML (e.g., WLS, DWLS)—are also a natural next step, given the generality of the M-estimation framework underlying RBM. Further investigation into non-normal likelihoods and robust estimation approaches would broaden the reach of RBM-based bias reduction to a wider class of psychometric and structural modelling problems.
@@ Appendix A Derivatives
is referred to as the structure matrix of $\mathbf{M}$ (petersen2012matrix, 100, Eq. 134,). Here, $\mathbf{J}^{ij}\in\mathbb{R}^{p\times q}$ is the single-entry matrix with a ‘$1$’ at the $(i,j)$ position and 0 elsewhere. Write
@@
using the fact that $d\tilde{\mathbf{B}}=\tilde{\mathbf{B}}(d\mathbf{B})\tilde{\mathbf{B}}$ (petersen2012matrix, 100, Eq. 59). Furthermore,